### v1

@@ 3.2.3 How does Census Town classification affect statutory recognition?
Although Census Town (CT) classification and statutory recognition are distinct processes that are built on different criteria and governed by different authorities, there is a strong relationship between the two and one often inspires the other (NIUA2020InterStateVar; roy2018predicting). Meeting the CT criteria signals that a settlement has urban characteristics and may warrant municipal governance. When a settlement crosses the CT thresholds, it draws attention from state urban development authorities and policymakers, who may then consider it for notification as a statutory town. The CT designation provides an objective benchmark indicating that the settlement has reached a level of population size, density, and economic activity consistent with urban areas. This can prompt state governments to evaluate whether the settlement is ready for municipal governance and whether it can sustain the administrative and fiscal responsibilities of a ULB. In fact, the central government also monitors CTs and encourages states to municipalize them, as exhorted by mohua2016 where the Ministry of Housing and Urban Affairs (MoHUA) formally urged states to consider CTs for statutory recognition.
@@ 3.3 Data Sources and Construction
Using SHRUG data platform444version 2.1 was available at the time of our study., we collect data at the village and town level [^5] This is the most granular level at which information is available and is identified using SHRUG IDs. for our study (asher2019shrug). We also collect state-level town directories from the Government of India’s Open Government Data (OGD) platform and list of Statutory Towns (STs) from National Housing Bank’s (NHB) website. We highlight the main datasets used in our analysis:
  • Population Census Data (2001 & 2011): We draw on comprehensive census data for all Indian settlements across two decennial rounds. The Primary Census Abstract (PCA) provides basic demographic characteristics including total population, gender and age composition, Scheduled Caste and Scheduled Tribe populations, literacy rates, and detailed workforce composition. These data allow us to construct the three Census Town eligibility criteria: population size, population density, and share of male main workers in non-agricultural activities. Further, we use the two village and town directory files on SHRUG -
    • Village Directory (Rural Areas): For rural settlements, we obtain detailed infrastructure and amenity data including area measures, education facilities at all levels (primary through senior secondary, both government and private), health facilities (hospitals, dispensaries, primary and community health centers, maternity centers, family welfare centers, tuberculosis clinics, veterinary hospitals), utilities (power supply availability, water supply sources, sanitation infrastructure), financial access (private banks, cooperative banks, agricultural credit societies), communication and transport infrastructure (post offices, telegraph offices, bus services, railways, paved roads), recreational facilities (public libraries, reading rooms, cinema halls, sports infrastructure), and distances to nearest urban facilities.
    • Town Directory (Urban Areas): For urban settlements, we collect infrastructure data analogous to the Village Directory (VD) with some differences in format and definition. These include area measures (in square kilometers), household counts, education facilities by level and type, health facilities with medical personnel counts, utilities differentiated by connection type, banking facilities, distances to key amenities when not present within town limits, and recreational and sports infrastructure.
  • Urban-Rural Classification: We use SHRUG’s official urban-rural classification identifiers to define each settlement’s status for the 2001 and 2011 census rounds. These classification keys link administrative identifiers (state and district codes) to unique SHRUG settlement identifiers, enabling us to match demographic and infrastructure data across datasets and accurately identify whether settlements are officially classified as urban or rural in each census year.
  • Statutory Town Lists: We obtain comprehensive lists of all statutorily recognized urban local bodies from the National Housing Bank and state-level census town directories from the Government of India’s Open Government Data (OGD) platform. These administrative records identify which settlements have been officially notified as municipalities, municipal corporations, town panchayats etc., regardless of whether they meet Census Town criteria. We merge these lists with census data using state, district, and settlement name identifiers to construct our treatment indicator for statutory recognition.
@@ 3.4 Running Variables and Treatment Assignment
We then define the treatment variable for statutory recognition based on whether a settlement was officially designated as a statutory town (ST) in 2011. This is a binary indicator that takes the value of 1 if the settlement was recognized as an ST in 2011, and 0 otherwise. The treatment assignment is not deterministic based on the running variables alone, as some settlements that meet the Census Town criteria may not receive statutory recognition due to state-level policy decisions, while others that do not meet the criteria may be recognized due to historical or political reasons. Therefore, we use a fuzzy regression discontinuity design that exploits the discontinuities in the probability of receiving statutory recognition at the thresholds defined by the running variables to identify the causal effect of urban governance on local public goods provision.
3.5 Key Outcome Variables
Outcomes come in two forms: presence/status indicators (e.g., existence of a primary school), counts (e.g., number of dispensaries), and in TD, distances to nearest facility. Education outcomes cover primary, middle, secondary, senior secondary, colleges (by stream), and vocational institutions. Health outcomes include total hospitals and dispensaries and staffing proxies (doctors/paramedics where available) and specialized centers (PHC/PHSC/MCW/FWC/TB/veterinary). Power outcomes record availability by sector; water/sanitation capture source types and drainage/processing. Access includes banks, markets, communications, transport modes, and road types. Appendix Appendix reports a full crosswalk between VD and TD variable names.
3.6 Sample and Summary Statistics
@@ 3.6.1 Urban/Rural Classification vs Statutory Recognition
We define statutory recognition using the state-level files available on Indian government’s Open Government Data (OGD) platform [^7] https://data.gov.in/. These files report the list of statutory towns (municipalities, corporations, nagar panchayats, etc.) for each state as of 2011. We merge these files with the PCA data using state and settlement names to create an indicator for statutory recognition.
Table below shows the cross-tabulation of statutory recognition and whether a settlement meets the CT thresholds.
@@
Not Statutory 544,894 26,848
Statutory 119 3,633
Panel B: Density $\geq$ 400 per km²
@@
Not Statutory 566,548 5,194
Statutory 1,175 2,577
Table 1 provides a descriptive sense of this process by cross-tabulating 2011 Census data on threshold eligibility and actual statutory status. Several patterns stand out. First, many settlements that exceed each individual threshold remain unrecognized. For instance, over 26,800 settlements had populations above 5,000 yet were not statutory towns in 2011, compared to only 3,633 settlements that were both above 5,000 and officially urban (Panel A of Table 1). A similar discrepancy is seen for the density criterion (Panel B) and the non-agricultural workforce criterion (Panel C). Second, looking at the combined criteria (Panel D), we see that a total of 7,771 settlements met all three urban criteria as of 2011. However, only 2,577 of these were actually recognized as statutory towns, while 5,194 settlements met the criteria but remained governed as rural. Conversely, there were 1,175 statutory towns that did not meet all the criteria – these are typically older small towns or special cases that had municipal status historically despite falling short on one of the Census metrics. This highlights that satisfying the Census definition is not sufficient for urban governance; political implementation by states is crucial. It also underscores the quasi-random aspect of our design: among settlements around the thresholds, some will have been converted to ULBs (treatment) and others not (control), not purely based on merit but partly due to state-specific policies and timing.
Recognizing a settlement as an urban local body brings a change in governance structure. Upon becoming a ULB, the settlement is governed by an elected urban local government (municipality or town council) rather than a rural council. Indian ULBs, as per the 74th Constitutional Amendment, are entrusted with a range of local functions. These include providing core public goods and services such as primary education, basic health and sanitation, water supply, solid waste management, street lighting, roads and transportation, and regulation of land use and construction. They also have authority to collect certain taxes (like property taxes) and user fees, and they receive transfers from state and central governments. In theory, municipalities should have greater resources and autonomy to invest in local infrastructure compared to rural bodies, although in practice many ULBs face capacity and funding constraints. We are interested in whether a settlement’s “switch” from rural to urban governance (statutory recognition) translates into measurable differences in local development indicators. Our hypothesis is that municipalities, despite their challenges, may deliver better access to public services than the rural governance system, due to a combination of dedicated funding, political accountability, and administrative focus on urban needs.
@@ 3.6.2 Treatment Assignment and Treatment Status
Notes: “CT in 2001” indicates whether a settlement met all three Census Town criteria (population $\geq$ 5,000, density $\geq$ 400 per km², non-agricultural workers $\geq$ 75%) in 2001. “Statutory by 2011” indicates whether the settlement had statutory town status in 2011. Panel B restricts to settlements within the bandwidth around the CT thresholds.
Several patterns emerge from Table 2. In the full sample (Panel A), the vast majority of settlements (98.5%) neither met the CT criteria in 2001 nor received statutory recognition by 2011. Among the 5,765 settlements that met the CT thresholds in 2001, only 2,542 (44%) gained statutory status by 2011, while 3,223 (56%) remained under rural governance despite being functionally urban. This substantial gap between eligibility and recognition illustrates the discretionary nature of state municipalization policy. Conversely, 1,199 settlements that did not meet the CT criteria nevertheless received statutory recognition, typically reflecting historical municipal status or special circumstances.
@@ 4 Empirical Strategy
In this section, we lay out the identification strategy. The goal is to estimate the causal effect of statutory recognition (municipal governance) on settlement outcomes.
@@ 4.1 Quasi-random variation from Census Towns
In this section, we show that Census Town (CT) designation leads to a discontinuous increase in the probability of statutory recognition for each of the thresholds. Furthermore, becoming a CT does not directly change local governance as the decision to grant a ULB is made by the state government. So theoretically, we should not expect a direct effect of CT designation on local outcomes coming from the local governance channel unless the settlement is also statutorily recognized. One could argue that CT may be capturing the differences in agglomeration economies or other unobserved characteristics that may be correlated with local outcomes. However, close to the thresholds, these characteristics should be similar on either side of the cutoff. For a country of 1.5 billion people, a settlement with a population of 4,900 and 5,100, or density of 390 and 410, is not likely to be systematically different in terms of agglomeration economies. Furthermore, even though we cannot rule out differences in unobserved characteristics, we can test for balance in observed characteristics (e.g., literacy rate, caste composition) and run placebo tests at pseudo-thresholds to check if there are any discontinuities in outcomes at these points. First, we show that there is a strong first stage relationship between meeting the CT thresholds and statutory recognition. We then use this quasi-random variation to instrument for statutory recognition and identify the Local Average Treatment Effect (LATE) of statutory recognition on local outcomes.
@@
$Z_{P,i}$ $=1(\frac{P_{i}}{5000}-1)\geq 0$
$Z_{D,i}$ $=1(\frac{D_{i}}{400}-1)\geq 0$
$Z_{N,i}$ $=1(\frac{N_{i}}{0.75}-1)\geq 0$
@@
Figure 1 presents graphically the first stage of our fuzzy RD design explained in equation 1 close to the thresholds [^8] We present results upto 100% away from the thresholds. The figure shows binned scatter plots of the probability of statutory status in 2011 against each running variable in 2001. Sub-figure (A) of Figure 1 shows the relationship between population and statutory recognition. Probability of statutory recognition is positively related to population, which is expected since an increase in population signifies a larger settlement that is more likely to be recognized as a town. Despite this positive relationship, there is a clear discontinuity in the probability of statutory recognition at the population threshold of 0 after normalization. This discontinuity indicates that settlements that just meet the population threshold are significantly more likely to receive statutory recognition than those that just miss it. Sub-figure (B) of Figure 1 shows the relationship between density and statutory recognition. Similar to population, there is a positive relationship between density and statutory recognition, but the discontinuity at the density threshold, while present, is less pronounced than for population. Sub-figure (C) of Figure 1 shows the relationship between non-agricultural share and statutory recognition. The relationship is flat, but there is a clear discontinuity at the non-agricultural share threshold. Settlements that just meet the non-agricultural share threshold are more likely to receive statutory recognition than those that just miss it. Overall, these figures provide evidence of a strong first stage relationship between meeting the CT thresholds and statutory recognition, which supports our identification strategy for estimating the causal effect of statutory recognition on local outcomes. In Section 5.1, we present a table that reports the first-stage coefficient and F-statistics for global and local estimates.
@@ 4.2 Fuzzy RD with multiple running variables
Figure 2: Probability of Treatment by Frontier Distance
Figure 2 shows the binned scatter plot of the probability of statutory recognition in 2011 against the minimum distance to any of the three thresholds (frontier distance) in 2001. As presented in the figure, there is a clear discontinuity in the probability of statutory recognition at the cutoff (frontier distance = 0). We implement a fuzzy RD design using this frontier distance as the running variable. Specifically, we estimate the following first stage and reduced form equations using observations close to the cutoff.
@@ 5.1 First stage: Effects on Statutory Recognition
Dependent variable: (Statutory town =1=1)
Instrument: CT eligibility in 2001 0.431∗∗∗ 0.071∗∗∗
@@
The first-stage results in Table 3 reveal important patterns regarding the strength and validity of our instrumental variable strategy. In the global sample, the instrument exhibits substantial explanatory power with a coefficient of 0.427 (standard error 0.0159) and a first-stage F-statistic of 717.68, well exceeding conventional thresholds for weak instrument concerns. However, this strong statistical relationship may be misleading for causal identification purposes. The global specification, while statistically powerful, is potentially vulnerable to endogeneity. Settlements across the entire population and density spectrum may differ systematically in ways that correlate with both census threshold compliance and subsequent outcomes. For instance, larger settlements may have greater political influence or administrative capacity that simultaneously makes them more likely to meet census criteria and more likely to attract infrastructure investment, regardless of statutory status. The high explanatory power (adjusted $R^{2}$ of 0.3233) suggests that the instrument is capturing substantial systematic variation that may not be quasi-random.
In contrast, the local specification of restricting analysis to settlements near the census cutoffs [^9] We use the following restrictions around the thresholds: (i) population $\pm 5000$ (ii) density $\pm 400$ (iii) non-ag male main work share $\pm 0.2$. provides more credible identification despite the smaller coefficient magnitude (0.071) and relatively lower F-statistic (18.81). This specification focuses on settlements that are observationally similar except for small differences in census characteristics that push them just above or below the thresholds. The lower adjusted $R^{2}$ (0.0581) and partial $R^{2}$ of instrument (0.0130) indicate that we are capturing variation that is more plausibly exogenous, as settlements just above and below the cutoffs should be similar in unobservable characteristics in the absence of treatment, as long as no manipulation occurs.
The F-statistic of 18.81 for our local specification is above the conventional threshold of 16.38 for strong instruments in the just-identified case (stock1997testing) and well above the critical value of 8.96 for 15% maximal IV bias (stock2002testing), suggesting the instrument maintains reasonable strength for our identification strategy. The local estimates thus provide our preferred specification for downstream analysis, offering a more credible foundation for interpreting the causal effects of statutory recognition on settlement outcomes.

### v2

@@ 2.2.2 How does Census Town classification affect statutory recognition?
Although Census Town (CT) classification and statutory recognition are distinct processes that are built on different criteria and governed by different authorities, there is a strong relationship between the two and one often inspires the other (Tewari, 2020; Roy and Pradhan, 2018). Meeting the CT criteria signals that a settlement has urban characteristics and may warrant municipal governance. When a settlement crosses the CT thresholds, it draws attention from state urban development authorities and policymakers, who may then consider it for notification as a statutory town. The CT designation provides an objective benchmark indicating that the settlement has reached a level of population size, density, and economic activity consistent with urban areas. This can prompt state governments to evaluate whether the settlement is ready for municipal governance and whether it can sustain the administrative and fiscal responsibilities of a ULB. In fact, the central government also monitors CTs and encourages states to municipalize them, as exhorted by Ministry of Housing and Urban Affairs, Government of India (2016) where the Ministry of Housing and Urban Affairs (MoHUA) formally urged states to consider CTs for statutory recognition.
@@ 2.3 Data Sources and Construction
Using SHRUG data platform444version 2.1 was available at the time of our study., we collect data at the village and town level [^5] This is the most granular level at which information is available and is identified using SHRUG IDs. for our study (Asher et al., 2019). We also collect state-level town directories from the Government of India’s Open Government Data (OGD) platform and list of Statutory Towns (STs) from National Housing Bank’s (NHB) website. We highlight the main datasets used in our analysis:
  • Population Census Data (2001 & 2011): We draw on comprehensive census data for all Indian settlements across two decennial rounds. The Primary Census Abstract (PCA) provides basic demographic characteristics including total population, gender and age composition, Scheduled Caste and Scheduled Tribe populations, literacy rates, and detailed workforce composition. These data allow us to construct the three Census Town eligibility criteria: population size, population density, and share of male main workers in non-agricultural activities.
  • Urban-Rural Classification: We use SHRUG’s official urban-rural classification identifiers to define each settlement’s status for the 2001 and 2011 census rounds.
  • Statutory Town Lists: We obtain comprehensive lists of all statutorily recognized urban local bodies from the National Housing Bank and state-level census town directories from the Government of India’s Open Government Data (OGD) platform.
@@ 2.4 Running Variables and Treatment Assignment
We then define the treatment variable for statutory recognition based on whether a settlement was officially designated as a statutory town (ST) in 2011. This is a binary indicator that takes the value of 1 if the settlement was recognized as an ST in 2011, and 0 otherwise. The treatment assignment is not deterministic based on the running variables alone, as some settlements that meet the Census Town criteria may not receive statutory recognition due to state-level policy decisions, while others that do not meet the criteria may be recognized due to historical or political reasons. Therefore, we use a fuzzy regression discontinuity design that exploits the discontinuities in the probability of receiving statutory recognition at the thresholds defined by the running variables to identify the causal effect of local urban governance status on local public goods provision.
2.5 Summary Statistics
@@ 2.5.1 Urban/Rural Classification vs Statutory Recognition
We define statutory recognition using the state-level files available on Indian government’s Open Government Data (OGD) platform [^7] State-level files were collected via https://data.gov.in/.. These files report the list of statutory towns (municipalities, corporations, nagar panchayats, etc.) for each state as of 2011. We merge these files with the PCA data using state and settlement names to create an indicator for statutory recognition.
@@
Not Statutory 485,580 17,509
Statutory 159 3,570
Panel B: Density $\geq$ 400 per km2
@@
Not Statutory 499,708 3,381
Statutory 1,194 2,535
Table 1 provides a descriptive sense of this process by cross-tabulating threshold eligibility in 2001 and statutory status by 2011. Several patterns stand out. First, 17,509 settlements had populations above 5,000 yet were not statutory towns by 2011, compared to 3,570 settlements that were both above 5,000 and statutorily recognized (Panel A of Table 1). A similar discrepancy is seen for the density criterion (Panel B), where 194,146 settlements had density above 400 per km2 in 2001 but were not statutory towns by 2011, while only 3,560 settlements met the density criterion and had statutory town status. For the non-agricultural workforce criterion (Panel C), 35,855 settlements had at least 75% non-agricultural male workers but were not statutory towns, compared to 2,673 that met this criterion in 2001 and were statutory towns by 2011. Second, looking at the combined criteria (Panel D), a total of 5,916 settlements met all three urban criteria in 2001 (3,381 not statutory + 2,535 statutory). However, only 2,535 of these were actually recognized as statutory towns by 2011, while 3,381 settlements met the criteria but remained governed as rural. Conversely, there were 1,194 statutory towns that did not meet all the criteria; these are typically older small towns or special cases that had municipal status historically despite falling short on one of the Census metrics. This highlights that satisfying the Census definition is not sufficient for urban governance; political implementation by states is crucial. It also underscores the quasi-random aspect of our design: among settlements around the thresholds, some received statutory recognition by 2011 and others did not, not purely based on merit but partly due to state-specific policies and timing. Nevertheless, as we will show in Section 3, crossing the thresholds greatly increases the likelihood of statutory recognition, providing the basis for our fuzzy regression discontinuity approach.
@@ 2.5.2 Treatment Assignment and Treatment Status
Notes: “CT in 2001” indicates whether a settlement met all three Census Town criteria (population $\geq$ 5,000, density $\geq$ 400 per km2, non-agricultural workers $\geq$ 75%) in 2001. “Statutory by 2011” indicates whether the settlement had statutory town status in 2011. Panel B restricts to settlements within the bandwidth around the CT thresholds.
Several patterns emerge from Table 2. In the full sample (Panel A), the vast majority of settlements (98.6%) neither met the CT criteria in 2001 nor received statutory recognition by 2011. Among the 5,765 settlements that met the CT thresholds in 2001, only 2,542 (44%) gained statutory status by 2011, while 3,223 (56%) remained under rural governance despite being functionally urban. This substantial gap between eligibility and recognition illustrates the discretionary nature of state municipalization policy. Conversely, 1,199 settlements that did not meet the CT criteria nevertheless received statutory recognition, typically reflecting historical municipal status or special circumstances.
@@ 3 Empirical Strategy
In this section, we lay out the identification strategy. The goal is to estimate the causal effect of statutory recognition and ULB status on settlement outcomes.
@@ 3.1 Quasi-random variation from Census Towns
In this section, we show that Census Town (CT) designation leads to a discontinuous increase in the probability of statutory recognition at the thresholds. Furthermore, becoming a CT does not directly change governance status as the decision to grant a ULB is made by the state government. So theoretically, we should not expect a direct effect of CT designation on local outcomes through the local urban governance status channel unless the settlement is also statutorily recognized. One could argue that CT may be capturing the differences in agglomeration economies or other unobserved characteristics that may be correlated with local outcomes. However, close to the thresholds, these characteristics should be similar on either side of the cutoff. For a country of 1.5 billion people, a settlement with a population of 4,900 and 5,100, or density of 390 and 410, is not likely to be systematically different in terms of agglomeration economies. Furthermore, even though we cannot rule out differences in unobserved characteristics, we can test for balance in observed characteristics (e.g., literacy rate, caste composition) and run placebo tests at pseudo-thresholds to check if there are any discontinuities in outcomes at these points. First, we show that there is a strong first-stage relationship between meeting the CT thresholds and statutory recognition. We then use this quasi-random variation to instrument for statutory recognition and identify the effect of statutory recognition on local outcomes.
@@
$Z_{P,i}$ $=\mathds{1}\!\left\{\frac{P_{i}}{5000}-1\geq 0\right\}$
$Z_{D,i}$ $=\mathds{1}\!\left\{\frac{D_{i}}{400}-1\geq 0\right\}$
$Z_{N,i}$ $=\mathds{1}\!\left\{\frac{N_{i}}{0.75}-1\geq 0\right\}$
@@
Notes: Each panel shows a binned scatter plot of the predicted probability of statutory recognition in 2011 against one of the three normalized running variables from 2001. Predicted probabilities are fitted values from the global first-stage OLS regression of statutory status on Census Town eligibility, controls, and district fixed effects (Equation 1). The dashed red line marks the normalized threshold.
Figure 1 presents graphically the first stage of our fuzzy RD design explained in equation 1 close to the thresholds [^8] We present results up to 100% away from the thresholds.. The figure shows binned scatter plots of the probability of statutory status in 2011 against each running variable in 2001. Sub-figure (A) of Figure 1 shows the relationship between population and statutory recognition. Probability of statutory recognition is positively related to population, which is expected since an increase in population signifies a larger settlement that is more likely to be recognized as a town. Despite this positive relationship, there is a clear discontinuity in the probability of statutory recognition at the population threshold of 0 after normalization. This discontinuity indicates that settlements that just meet the population threshold are significantly more likely to receive statutory recognition than those that just miss it. Sub-figure (B) of Figure 1 shows the relationship between density and statutory recognition. Similar to population, there is a positive relationship between density and statutory recognition, but the discontinuity at the density threshold, while present, is less pronounced than for population. Sub-figure (C) of Figure 1 shows the relationship between non-agricultural share and statutory recognition. The relationship is flat, but there is a clear discontinuity at the non-agricultural share threshold. Settlements that just meet the non-agricultural share threshold are more likely to receive statutory recognition than those that just miss it. Overall, these figures provide evidence of a strong first-stage relationship between meeting the CT thresholds and statutory recognition, which supports our identification strategy for estimating the causal effect of statutory recognition on local outcomes. In Section 4.1, we present a table that reports the first-stage coefficient and F-statistics for global and local estimates.
@@ 3.2 Fuzzy RD with multiple running variables
Figure 2: Probability of Statutory Recognition by Frontier Distance
Notes: Binned scatter plot of the predicted probability of statutory recognition in 2011 against the soft-min frontier running variable constructed from the three 2001 Census Town criteria. The frontier variable is defined as $r^{*}_{i}=-\frac{1}{\alpha}\ln\bigl(\exp(-\alpha\,r_{P})+\exp(-\alpha\,r_{D})+\exp(-\alpha\,r_{N})\bigr)$ with $\alpha=10$, which smoothly approximates the minimum of the three normalized distances. Negative values indicate that a settlement falls short of at least one threshold; positive values indicate that all three thresholds are met. Each dot represents the mean predicted probability within an equal-width bin. Predicted probabilities are fitted values from the global first-stage regression [^9] For our main analysis in Section 4.2, we use local first-stage estimates. (Equation 1).
Figure 2 plots the predicted probability of statutory recognition against the frontier running variable, which collapses the three Census Town dimensions into a single scalar. The figure reveals a clear discontinuity at the eligibility frontier, near the zero on the horizontal axis. To the left of the cutoff, where settlements fall short of at least one threshold, the predicted probability of receiving a municipal body is close to zero and hovers near 1%. Immediately to the right, where settlements satisfy all three criteria, the probability jumps sharply to approximately 35%, rising further to above 40% for settlements that comfortably exceed the thresholds. This pattern is precisely what the fuzzy RD design requires: a large, discrete change in the probability of treatment at the eligibility boundary, embedded within an otherwise smooth relationship. The magnitude of the jump is substantial and visually unambiguous, providing strong graphical evidence that Census Town eligibility is a predictor of statutory recognition. This reinforces the relevance condition of our instrument and is consistent with the formal first-stage F-statistics reported in Table 3.
The contrast between Figure 2 and the individual-threshold plots in Figure 1 is also informative. When projected onto any single running variable, the discontinuity appears modest because most of the variation in eligibility is driven by the joint interaction of all three criteria. The frontier variable captures this joint variation and motivates our use of a multi-dimensional instrument defined by the intersection of all three thresholds, rather than relying on any single cutoff.
@@ 4.1 First stage: Effects on Statutory Recognition
Dependent variable: Statutory town =1=1
CT eligibility in 2001 0.431∗∗∗ 0.071∗∗∗
@@
The first-stage results in Table 3 reveal important patterns regarding the strength and validity of our instrumental variable strategy. In the global sample, the instrument exhibits substantial explanatory power with a coefficient of 0.431 and a first-stage F-statistic of 749.43, well exceeding conventional thresholds for weak instrument concerns. However, this strong statistical relationship may be misleading for causal identification purposes. The global specification, while statistically powerful, is potentially vulnerable to endogeneity. Settlements across the entire population and density spectrum may differ systematically in ways that correlate with both census threshold compliance and subsequent outcomes. For instance, larger settlements may have greater political influence or administrative capacity that simultaneously makes them more likely to meet census criteria and more likely to attract infrastructure investment, regardless of statutory status.
In contrast, the local specification of restricting analysis to settlements near the census cutoffs [^10] We use the following restrictions around the thresholds: (i) population $\pm 5000$ (ii) density $\pm 400$ (iii) non-ag male main work share $\pm 0.2$. Table 11 in the appendix presents bandwidth-sensitivity results. provides more credible identification despite the smaller coefficient magnitude (0.071) and relatively lower F-statistic (18.05). This specification focuses on settlements that are observationally similar except for small differences in census characteristics that push them just above or below the thresholds. The lower adjusted $R^{2}$ (0.1007) and partial $R^{2}$ of instrument (0.0127) indicate that we are capturing variation that is more plausibly exogenous, as settlements just above and below the cutoffs should be similar in unobservable characteristics in the absence of treatment, as long as no manipulation occurs.
The F-statistic of 18.05 for our local specification is above the conventional threshold of 16.38 for strong instruments in the just-identified case (Staiger and Stock, 1997) and well above the critical value of 8.96 for 15% maximal IV bias (Stock and Yogo, 2002), suggesting the instrument maintains reasonable strength for our identification strategy. The local estimates thus provide our preferred specification for downstream analysis, offering a credible foundation for interpreting the causal effects of statutory recognition on settlement outcomes.